OpenStudy (anonymous):

Write an equation that satisfies the condition: slope: 4/5, passes through (10,-3). Please show work if possible thanks!

OpenStudy (anonymous):

y=mx+b is the slope intercept form Sub in the point and the slope -3=4/5(10)+b -3=8+b -11=b equation is y=4/5x-11

OpenStudy (anonymous):

\[y-y _{1}=m(x-x _{1})\] Just plug in your point and your slope and manipulate it to the form you need.

OpenStudy (anonymous):

thank you both!

OpenStudy (anonymous):

Or what TBates said, that works too.